Please show all steps.
Prove that if f1(n) = N2(91(n)) and f(n) = 12(92(n)), then f1(n) + f2(n) = N(91(n) +g2 (n))
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

By the definition of Big Omega

f1(n)=Omega(g1(n))

So,

f1(n)>=c1*g(n)

f2(n)=Omega(g2(n))

So,

f2(n)>=c2*g2(n)

Add both

So,

f1(n)+f2(n)>=(c1*g1(n)+c2*g2(n))>=max(c1,c2)*(g1(n)+g2(n))

So,

We found constant c=max(c1,c2)

So, by definition

f1(n)+f2(n)=Omega(g1(n)+g2(n))

Kindly revert for any queries

Thanks.